Graph the given​ functions, f and​ g, in the same rectangular coordinate system. Describe how the graph of g is related to the graph of
f(x)= -x^3
​g(x)= -x^3-2
1
Expert's answer
2020-08-24T18:49:17-0400
"Solution"

Given

"f(x)= -x^3,\\\\\n\n\u200bg(x)= -x^3-2"


Drawing the graph of the given functions and finding the relationship between them


Graph -



Relation -


g(x) is related to f(x) in that it is the same graph but shifted down by 2 units.